A 9.00-kg hanging object is connected by a light,inextensible cord over a light ,frictionless pulley to a 5.00-kg block that is sliding on a flat table. Taking the coefficient of kinetic friction as -.200, find the tension in the string


1
Expert's answer
2021-09-13T15:44:06-0400


the arrow in brown shows the friction,

block m2:

"m_2 g - T = m_2 a \\

9 \times9.8 - T = 9a \;\;\; (1) \\

N - m_2 g = 0 \\

N = m2 g = 5 \times 9.8 = 49 \; N \\

f = μN \\

f = 0.2 \times 49 = 9.8 \;N"

block m1:

"T - f = m_1a \\

T - 9.8 = 5a \;\;\;(2)"

solving 1 and 2:

"T = 37.8 \; N \\


a = 5.6 \;m/s^2"